a quantity of low density gas in rigid container is initially at room temperature ( 20 degrees C ) and a presure of 15 kpa. what will be the pressure if the temperature increase thrice its initial temperature?


1
Expert's answer
2021-10-18T11:11:15-0400

The Charles's law (constant volume processes) says

"\frac{P_1}{T_1}=\frac{P_2}{T_2}"

Hence, the final temperature

"T_2=\frac{P_2}{P_1}T_1=\frac{3P_1}{P_1}(20+273)=879\:\rm K=606^{\circ}C"